OpenStudy (anonymous):

Using the Pythagorean theorem, find the measure of the hypotenuse when a = 12mm and b = 25mm. A. c = 26mm B. c = 26.7mm C. c = 27mm D. c = 27.7mm

OpenStudy (oaktree):

12^2+25^2=144+625=769, so the answer is sqrt(769) = 27.7.

OpenStudy (anonymous):

Thanks (: